OpenStudy (ellamoyseyuk):

Jack and Maureen are opening an ice hotel. The structure and furniture are made out of ice. They must determine the weight of some of the ice furniture to decide the size of the crane needed to lift each piece into place. Maureen knows that ice weighs approximately 62.4 pounds per cubic foot. Find the approximate weight of each of the figures below. a. The stage in the banquet hall measures 22 ft by 15 ft by 4 ft. b. Two cylindrical pillars will flank the front door. Each cylinder has diameter 1 ft and height 9 ft. c. A series of four spheres will decorate the reception area. The largest sphere has diameter 4 ft.The other three spheres have diameter 2 ft, 1 ft, and 1/2 ft.

OpenStudy (prathamesh_m):

Find the volume of each of those structures in cubic feet. Then to calculate the weight multiply the volume by 62.4 since one cubic foot weighs 62.4 pounds.

OpenStudy (prathamesh_m):

Option a: The stage is a cuboid, its volume is obtained by multiplying its three dimensions(length x breadth x height) Option b: The volume of one cuboid is given by: \[V = 2\pi rh\] where r is radius and h is height of the pillar Your third option is incomplete.

OpenStudy (ellamoyseyuk):

a) 22*15*4 = 1320 ft.volume = 1320*62.4 = 82368 b) v = 2 pi r*h, v = 2pi*1/2*9 = 28.2743338823 + 28.2743338823 =56.5486677646 ft. volume = 56.5486677646*62.4 = 3528.63686851 ft

OpenStudy (prathamesh_m):

Yes you're right.

OpenStudy (ellamoyseyuk):

A) V = l*w*h V = 22*15*4 V = 1320 = 1320*62.4 = 82368 pounds. B) V = Bh V = pi ½^2*9 V = pi*0.25*9 V = 2.25 pi V = 7.07 = 7.07 * 62.4 = 441.168 pounds. V = Bh V = pi ½^2*9 V = pi*0.25*9 V = 2.25 pi V = 7.07 = 7.07 *62.4 = 441.168 pounds C) v= 4/3pir^3 V = 4/3pi*4^3 v =4/3 pi*64 v = 85.33 pi v = 268.07 = 268.07 * 62.4 = 16727.568 pounds v= 4/3pir^3 V = 4/3pi*2^3 v =4/3 pi*8 v = 10.67 pi v = 33.52 = 33.52 * 62.4 = 2091.648 pounds. v= 4/3pir^3 V = 4/3pi*1^3 v =4/3 pi*1 v = 1.33 pi v = 4.18 = 4.18 * 62.4 = 260.832 pounds. v= 4/3pir^3 V = 4/3pi*1/2^3 v =4/3 pi*0.125 v = 0.17 pi v = 0.53 = 0.53 * 62.4 = 33.072 pounds.